consider a bond with a face value of $100 paying an annual coupon of $20 and maturing in two years. the one-year interest rate is 10% and the two-year interest rate is 7%. what is the yield to maturity of the bond?
Answer #1

one-year interest rate = z1 = 10%

two-year interest rate = z2 = 7%

Annual coupon, C = 20

Face value, F = 100

Price of the bond, P = C / (1 + z1) + (F + C) / (1 + z2)2 = 20 / (1 + 10%) + (100 + 20) / (1 + 7%)2 =  122.99

Yield to maturity (YTM) can now be calculated using the RATE function of excel. Inputs are:

  • Nper = 2 years
  • PMT = C = 20
  • PV = - Price of the bond = - 122.99
  • FV = F = 100

Hence, YTM = RATE (Nper, PMT, PV, FV) = RATE (2, 20, -122.99, 100) = 7.24%